OpenStudy (anonymous):

Tommy can paint a fence in 12 hours. If his friends help they can paint the fence in

OpenStudy (anonymous):

1 hour.

OpenStudy (anonymous):

How long would it take for huck to paint he fence by himself

OpenStudy (anonymous):

guessing 12 hours

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

16.4 hours

OpenStudy (anonymous):

Let z be the time that his friend can paint the fence by himself than in one ha to solve the equation \[ \frac{1}{z}+\frac{1}{12}=\frac 11 \] why?

OpenStudy (anonymous):

\[ 12 + z =12 z\\ 11 z =12\\ z= \frac {11}{12} \]

OpenStudy (anonymous):

Practice problems related to your on http://www.saab.org/mathdrills/workcompletion.cgi